The Iarussi Legal Aide office assisted 126 people in June 2013, with a staff labor cost of $3,240. In June 2014, the office provided assistance to 145 people with a labor cost of $3,960.

The productivity growth over this one-year period is ______% customers assisted per labor $. (If your answer is negative value, please enter "minus" sign.)
Answer #1

Productivity year 2013 = 126 / 3240 = 0.0388

Productivity year 2014 = 145 / 3960 = 0.0366

Change in productivity =( (0.0366 - 0.0388) / 0.0388) * 100% = -5.67 %